OpenStudy (anonymous):

3x-4y=3 I have to solve the equation for y

OpenStudy (anonymous):

yeah possibly

OpenStudy (anonymous):

to solve for x replace y for zero. to solve for y replace x with zero

OpenStudy (anonymous):

Y = 3(0)-4y=3

OpenStudy (anonymous):

X = 3x-4(0)=3

OpenStudy (anonymous):

To solve for Y use -4y=3

OpenStudy (anonymous):

y=-3/4

OpenStudy (anonymous):

OR -.75

OpenStudy (anonymous):

Sodo I subtract 3x from both sides or add 3x to both sides??

OpenStudy (anonymous):

you divide 3 from both sides to isolate the variable